OpenStudy (anonymous):

Match each form with its equation. Answer Slope-intercept form Read Answer Items for Question 1 Point-slope form Read Answer Items for Question 1 Standard form Read Answer Items for Question 1 Horizontal Line Read Answer Items for Question 1 Vertical Line Read Answer Items for Question 1 Answer A. x = c B. y = b C. y = mx + b D. Ax + By = C E. y - y1 = m(x - x1)

OpenStudy (anonymous):

Slope Intercept: y=mx +b Point Slope: y - y1 = m(x - x1) Standard Form: Ax + By = C Horizontal Line y= b Vertical Line: x =c

OpenStudy (anonymous):

1. C 2. E 3. D 4. B 5. A

OpenStudy (anonymous):

Thanks...

OpenStudy (anonymous):

can you help me with the other questions?